# A probabilistic, data-driven closure model for RANS simulations with aleatoric, model uncertainty

### v1

@@ 2.2 Probabilistic, model-consistent data-driven differential framework
where $\bm{z}=[\bm{u},\bm{p}]^{T}$ summarily denotes the discretized velocity $\bm{u}$ and pressure $p$ fields and $\bm{\tau}$ the discretized RS field. E.g. for a two-dimensional flow domain $\bm{z}\in\mathbb{R}^{N\times 3}$, $\bm{\tau}\in\mathbb{R}^{N\times 3}$ where $N$ is the number of grid points. The discretization scheme employed and other implementation details are discussed in A. We denote with $\mathcal{G}$ the discretized, non-linear operator accounting for the advective and diffusive terms on the left-hand side of Equation (7) as well as the conservation of mass in Equation (8), and with $\bm{B}$ the matrix (i.e. linear operator) arising from the divergence term on the right-hand side of Equation (7).
@@
Traditional, data-driven strategies postulate a closure e.g. $\bm{\tau}_{\bm{\theta}}(\bm{z})$ (or most often $\bm{\tau}_{\bm{\theta}}(\bm{u})$) dependent on some tunable parameters $\bm{\theta}$, which they determine either by assuming that reference Reynolds-stress data is available from DNS simulations (or in general, from higher-fidelity models such as LES) or by employing experimental or simulation-based data of the mean velocities/pressures i.e. of $\bm{z}$. The former scenario which is referred to as model regression Ahmed et al. (2021) has received significant attention in the past (e.q. Geneva and Zabaras (2019); Ling et al. (2016b); Wang et al. (2017b); Kaandorp and Dwight (2020)). Apart from the heavier data requirements, it does not guarantee that the trained model would yield accurate predictions of $\bm{z}$ Thompson et al. (2016) as even small errors in $\bm{\tau}$ might get amplified when solving Equation (12). The second setting, referred to as trajectory regression in Ahmed et al. (2021), might be able to make use of indirect and noisy observations but is much more cumbersome as repeated model evaluations and parametric sensitivities, i.e. a differentiable solver, are needed for training.
@@
Critical to any data-driven model is its ability to generalize i.e. to produce accurate predictions under different flow scenaria. On one hand this depends on the training data available but on the other on incorporating a priori available domain knowledge. The latter can attain various forms and certainly includes known invariances or equivariances that characterize the associated maps. Apart from this and the particulars of the parameterized model form, a critical aspect pertains to uncertainty quantification. We distinguish between parametric and model uncertainty. The former is of epistemic origin and has been extensively studied (e.g.,Oliver and Moser (2009, 2011); Edeling et al. (2014)). Bayesian formulations offer a rigorous manner for quantifying it and ultimately propagating it in the predictive estimates in the form of the predictive posterior. We note however that in the limit of infinite data, the posterior of the model parameters $\bm{\theta}$ (no matter what these are or represent) would collapse to a Dirac-delta i.e. point-estimates for $\bm{\theta}$ would be obtained. This false lack of uncertainty does not imply that the model employed is perfect as the true (unknown) closure might attain a form not contained in the parametric family used or in the features of $\bm{z}$ that appear in the input (e.g. even though all models proposed employ a locality assumption in the closure equations, non-local features of $\bm{u}$ might be needed).
@@
The issue of model uncertainty in the closure equations which is of aleatoric nature, has been much less studied and represents the main contribution of this work. In particular, we augment the parametric closure model $\bm{\tau}_{\bm{\theta}}(\bm{u})$ with a set of latent (i.e. unobserved) random variables $\bm{\epsilon}_{\tau}$ which are embedded in the model equations and which quantify model discrepancies at each grid point. In reference to the discretized RS vector $\bm{\tau}$ in Equation (12), we propose:
@@ 2.2.1 Parametric RS model
In this section we discuss the parametric part, i.e. $\bm{\tau}_{\bm{\theta}}(\bm{u})$ in the closure model of Equation (13). As this represents a vector containing its values at various grid points over the problem domain, the ensuing discussion and equations should be interpreted as per grid point. We note that the most popular LEVM model (Equation (10)) assumes that the anisotropic part of the $\bm{\tau}_{LEVM}$, is linearly related to the mean strain rate tensor $\bm{\bar{S}}$. This linear relation assumption restricts the model to attain a small subset of all the possible states of turbulence. This subset is referred to as the plane strain line Iaccarino et al. (2017). Experimental and DNS data show, that turbulent flows explore large regions of the domain of realizable turbulence states.
In the present work, we make use of the invariant neural network architecture proposed by Ling et al. (2016b) which relates the anisotropic part of the RS tensor with the symmetric and antisymmetric components of the velocity gradient tensor. By using tensor invariants, the neural network is able to achieve both Galilean invariance as well as rotational invariance. The Navier-Stokes equations are Galilean-invariant, i.e. they remain unchanged for all inertial frames of reference. The theoretical foundation of this neural network lies in the Non-Linear Eddy Viscosity Model (NLEVM) proposed by Pope (1975) and has been used in several studies Geneva and Zabaras (2019); Xiao et al. (2016); Kaandorp and Dwight (2020). By employing barycentric realizability maps Gorlé and Iaccarino (2013); Emory et al. (2013); Mishra and Iaccarino (2017); Iaccarino et al. (2017); Thompson et al. (2019), it was shown in Kaandorp and Dwight (2020) that this architecture overcomes the plane strain line restriction and can explore other realizable states.
In the model proposed by Pope (1975), the normalized anistropic tensor of the R-S was given by $\bm{b}:=\bm{b}(\bm{S},\bm{\Omega})$, which was a function of the normalized mean rate of strain tensor $\bm{S}$ and the rotation tensor $\bm{\Omega}$, i.e.:
@@
Through the application of Cayley-Hamilton theorem Pope (1975), the following general expression for the anisotropy tensor $\bm{b}$ was adopted:
$\bm{b}=\sum_{k=1}^{10}G^{(k)}(\underbrace{\mathcal{I}_{1}...\mathcal{I}_{5}}_{\begin{subarray}{c}Scalar\\ Invarients\end{subarray}})\bm{\mathcal{T}}^{(k)}$ (17)
@@
$\mathcal{I}_{1}=\mathrm{tr}(\bm{S}^{2}),\quad\mathcal{I}_{2}=\mathrm{tr}(\bm{\Omega}^{2}),\quad\mathcal{I}_{3}=\mathrm{tr}(\bm{S}^{3}),\quad\mathcal{I}_{4}=\mathrm{tr}(\bm{\Omega}^{2}\bm{S}),\quad\mathcal{I}_{5}=\mathrm{tr}(\bm{\Omega}^{2}\bm{S}^{2}).$ (18)
and $\mathcal{T}^{(k)}$ are the symmetric tensor basis functions (the complete set is listed in Table 1). The coefficients $G^{(i)}$ are scalar, non-linear functions which depend on the five invariants $\mathcal{I}_{1}...\mathcal{I}_{5}$ and must be determined. If $G^{(1)}=-0.09,G^{(n)}=0$, the NLEVM degenerates to the classical $k-\epsilon$ model. When the NLEVM was proposed, it was impossible to find good approximations for these functions and as a result it did not receive adequate attention. This hurdle however was overcome with the help of machine learning Ling et al. (2016b) where $G^{(i)}$ were learned from high-fidelity simulation data. Neural networks with parameters $\bm{\theta}$ were employed for the coefficients i.e. $G_{\bm{\theta}}^{(i)}$ and:
$\bm{b}_{\bm{\theta}}=\sum_{i=1}^{10}G_{\bm{\theta}}^{(i)}(\underbrace{\mathcal{I}_{1}...\mathcal{I}_{5}}_{\begin{subarray}{c}Scalar\\ Invarients\end{subarray}})\bm{\mathcal{T}}^{(i)};\quad\bm{\tau}_{\bm{\theta}}=2k\bm{b}_{\bm{\theta}}+\frac{2k}{3}\bm{I};$ (21)
@@
As in Geneva and Zabaras (2019), we employ the following prior for the NN parameters $\bm{\theta}$:
$p(\bm{\theta}\mid\nu)=\mathcal{N}(\bm{\theta}|0,\nu^{-1}\bm{I}_{d_{\theta}}),\quad p(\nu)=Gamma(\nu|a_{0},b_{0})$ (22)
where $d_{\theta}=dim(\bm{\theta})$ and a Gamma hyperprior was used for the common precision hyperparameter $\nu$ with $(a_{0},b_{0})=(1.0,0.02)$. The resulting prior has the density of a Student’s $\mathcal{T}$-distribution centered at zero, which is obtained by analytically marginalizing over the hyperparameter $\nu$ Bishop and Nasrabadi (2006).
@@
$\bm{\mathcal{T}}^{(1)}$
= $\bm{S}$,
$\bm{\mathcal{T}}^{(6)}$
= $\bm{\Omega}^{2}\bm{S}+\bm{S}\bm{\Omega}^{2}-\frac{2}{3}\mathrm{tr}(\bm{S}\bm{\Omega}^{2})\bm{I}$,
$\bm{\mathcal{T}}^{(2)}$
= $\bm{S}\bm{\Omega}-\bm{\Omega}\bm{S}$,
$\bm{\mathcal{T}}^{(7)}$
= $\bm{\Omega}\bm{S}\bm{\Omega}^{2}+\bm{\Omega}^{2}\bm{S}\bm{\Omega}$,
$\bm{\mathcal{T}}^{(3)}$
= $\bm{S}^{2}-\frac{1}{3}\mathrm{tr}(\bm{S}^{2})\bm{I}$,
$\bm{\mathcal{T}}^{(8)}$
= $\bm{S}\bm{\Omega}\bm{S}^{2}-\bm{S}^{2}\bm{\Omega}\bm{S}$,
$\bm{\mathcal{T}}^{(4)}$
= $\bm{\Omega}^{2}-\frac{1}{3}\mathrm{tr}(\bm{\Omega}^{2})\bm{I}$,
$\bm{\mathcal{T}}^{(9)}$
= $\bm{\Omega}^{2}\bm{S}^{2}+\bm{S}^{2}\bm{\Omega}^{2}-\frac{2}{3}\mathrm{tr}(\bm{S}^{2}\bm{\Omega}^{2})\bm{I}$,
$\bm{\mathcal{T}}^{(5)}$
= $\bm{\Omega}\bm{S}^{2}-\bm{S}^{2}\bm{\Omega}$,
$\bm{\mathcal{T}}^{(10)}$
= $\bm{\Omega}\bm{S}^{2}\bm{\Omega}^{2}-\bm{\Omega}^{2}\bm{S}^{2}\bm{\Omega}$,
@@ 2.2.2 Stochastic discrepancy term to RS
$\bm{\epsilon}_{\tau}=\bm{W}\bm{E_{\tau}}$ (23)
It is based on considering $N_{d}$ subdomains of the problem domain and assuming that for all grid points in a certain subdomain, the corresponding RS discrepancy terms are identical. This can be expressed as in Equation (23) above where the entries of $\bm{W}$ are $1$ if a corresponding grid point (row of $\bm{W}$) belongs in a certain subdomain (column of $\bm{W}$) and 00 otherwise. The vector $\bm{E}_{\tau}=\{\bm{E}_{\tau,J}\}_{J=1}^{N_{d}}$ contains therefore the RS discrepancy terms for each subdomain, e.g. $dim(\bm{E}_{\tau,J})=3$ for a two-dimensional flow. In the ensuing numerical illustrations, the division into subdomains is done in a regular manner and $\bm{W}$ is prescribed a priori. One could nevertheless readily envision a learnable $\bm{W}$ or even an adaptive refinement into subdomains.
The incorporation of the model discrepancy variables $\bm{\epsilon}_{\tau}$ or $\bm{E}_{\tau}$ at each grid-point or subdomain introduces redundancies i.e. there would be an infinity of combinations of $\bm{\theta}$ and $\bm{\epsilon}_{\tau}$/$\bm{E}_{\tau}$ that could fit the data equally well. In order to address this issue, we invoke the concept of sparsity which has been employed in similar situations in the context of physical modeling Brunton et al. (2016); Felsberger and Koutsourelakis (2019). To this end, we make use of a sparsity-enforcing Bayesian prior based on the Automatic Relevance Determination (ARD) Neal (2012); Wu et al. (2014). In particular:
$p(\bm{E}_{\tau}|\bm{\Lambda})=\prod_{J=1}^{N_{d}}p(\bm{E}_{\tau,J}|\bm{\Lambda}^{(J)})=\prod_{J=1}^{N_{d}}\mathcal{N}\left(\bm{E}_{\tau,J}~{}|\bm{0},~{}diag(\bm{\Lambda}_{J})^{-1})\right)$ (24)
@@
$p(\bm{\Lambda})=\prod_{J=1}^{N_{d}}\prod_{\ell=1}^{L}Gamma(\Lambda_{J,\ell}~{}|\alpha_{0},\beta_{0})$ (25)
where $\Lambda_{J,\ell}$ denotes the $\ell^{th}$ entry (e.g. $L=3$ for two-dimensional flows) of the vector of precision hyperparameters in subdomain $J$. We note that when $\Lambda_{J,\ell}\to\infty$, then the corresponding model discrepancy term $E_{\tau,J,\ell}\to 0$. The resulting prior for $\bm{E}_{\tau}$ arising by marginalizing the hyperparameters $\bm{\Lambda}$ is a light-tailed, Student’s t-distribution Tipping (2000) that promotes solutions in the vicinity of 00 unless strong evidence in the data suggests otherwise. The hyperparameters $\alpha_{0},\beta_{0}$ are effectively the only ones that need to be provided by the analyst. We advocate very small values ($\alpha_{0}=\beta_{0}=10^{-3}$ was used in the ensuing numerical illustrations) which correspond to a uninformative prior.

### v2

@@ 2.2 Probabilistic, model-consistent data-driven differential framework
where $\bm{z}=[\bm{u},{\color[rgb]{0,0,0}p}]^{T}$ summarily denotes the discretized velocity $\bm{u}$ and pressure $p$ fields and $\bm{\tau}$ the discretized RS field. E.g. for a two-dimensional flow domain $\bm{z}\in\mathbb{R}^{N\times 3}$, $\bm{\tau}\in\mathbb{R}^{N\times 3}$ where $N$ is the number of grid points. The discretization scheme employed and other implementation details are discussed in A. We denote with $\mathcal{G}$ the discretized, non-linear operator accounting for the advective and diffusive terms on the left-hand side of Equation (7) as well as the conservation of mass in Equation (8), and with $\bm{B}$ the matrix (i.e. linear operator) arising from the divergence term on the right-hand side of Equation (7).
@@
Traditional, data-driven strategies postulate a closure e.g. $\bm{\tau}_{\bm{\theta}}(\bm{z})$ (or most often $\bm{\tau}_{\bm{\theta}}(\bm{u})$) dependent on some tunable parameters $\bm{\theta}$, which they determine either by assuming that reference Reynolds-stress data is available from DNS simulations (or in general, from higher-fidelity models such as LES) or by employing experimental or simulation-based data of the mean velocities/pressures i.e. of $\bm{z}$. The former scenario which is referred to as model regression [64] has received significant attention in the past (e.q. [22, 15, 65, 17]). Apart from the heavier data requirements, it does not guarantee that the trained model would yield accurate predictions of $\bm{z}$ [24] as even small errors in $\bm{\tau}$ might get amplified when solving Equation (12). The second setting, referred to as trajectory regression in [64], might be able to make use of indirect and noisy observations but is much more cumbersome as repeated model evaluations and parametric sensitivities, i.e. a differentiable solver, are needed for training.
@@
Critical to any data-driven model is its ability to generalize i.e. to produce accurate predictions under different flow scenaria. On one hand this depends on the training data available but on the other, on incorporating a priori available domain knowledge. The latter can attain various forms and certainly includes known invariances or equivariances that characterize the associated maps. Apart from this and the particulars of the parameterized model form, a critical aspect pertains to uncertainty quantification. We distinguish between parametric and model uncertainty. The former is of epistemic origin and has been extensively studied (e.g., [9, 42, 43]). Bayesian formulations offer a rigorous manner for quantifying it and ultimately propagating it in the predictive estimates in the form of the predictive posterior. We note however that in the limit of infinite data, the posterior of the model parameters $\bm{\theta}$ (no matter what these are or represent) would collapse to a Dirac-delta i.e. point-estimates for $\bm{\theta}$ would be obtained. This false lack of uncertainty does not imply that the model employed is perfect as the true (unknown) closure might attain a form not contained in the parametric family used or in the features of $\bm{z}$ that appear in the input (e.g. even though all models proposed employ a locality assumption in the closure equations, non-local features of $\bm{u}$ might be needed).
@@
The issue of model uncertainty in the closure equations which is of an aleatoric nature, has been much less studied and represents the main contribution of this work. In particular, we augment the parametric closure model $\bm{\tau}_{\bm{\theta}}(\bm{u})$ with a set of latent (i.e. unobserved) random variables $\bm{\epsilon}_{\tau}$ which are embedded in the model equations and which quantify model discrepancies at each grid point. In reference to the discretized RS vector $\bm{\tau}$ in Equation (12), we propose:
@@ 2.2.1 Parametric RS model
In this section we discuss the parametric part, i.e. $\bm{\tau}_{\bm{\theta}}(\bm{u})$ in the closure model of Equation (13). As this represents a vector containing its values at various grid points over the problem domain, the ensuing discussion and equations should be interpreted as per grid point. We note that the most popular LEVM model (Equation (10)) assumes that the anisotropic part of the $\bm{\tau}_{LEVM}$, is linearly related to the mean strain rate tensor $\bm{\bar{S}}$. This linear relation assumption restricts the model to attain a small subset of all the possible states of turbulence. This subset is referred to as the plane strain line [66]. Experimental and DNS data show that turbulent flows explore large regions of the domain of realizable turbulence states.
In the present work, we make use of the invariant neural network architecture proposed by [15] which relates the anisotropic part of the RS tensor with the symmetric and antisymmetric components of the velocity gradient tensor. By using tensor invariants, the neural network is able to achieve both Galilean invariance as well as rotational invariance. The Navier-Stokes equations are Galilean-invariant, i.e. they remain unchanged for all inertial frames of reference. The theoretical foundation of this neural network lies in the Non-Linear Eddy Viscosity Model (NLEVM) proposed by [21] and has been used in several studies [22, 13, 17]. By employing barycentric realizability maps [46, 45, 67, 66, 48], it was shown in [17] that this architecture overcomes the plane strain line restriction and can explore other realizable states.
In the model proposed by [21], the normalized anisotropic tensor of the R-S was given by $\bm{b}:=\bm{b}(\bm{S},\bm{\Omega})$, which was a function of the normalized mean rate of strain tensor $\bm{S}$ and the rotation tensor $\bm{\Omega}$, i.e.:
@@
Through the application of Cayley-Hamilton theorem [21], the following general expression for the anisotropy tensor $\bm{b}$ was adopted:
$\bm{b}=\sum_{k=1}^{10}G^{(k)}(\underbrace{\mathcal{I}_{1}...\mathcal{I}_{5}}_{\begin{subarray}{c}Scalar\\ Invariants\end{subarray}})\bm{\mathcal{T}}^{(k)},$ (15)
@@
$\mathcal{I}_{1}=\mathrm{tr}(\bm{S}^{2}),\quad\mathcal{I}_{2}=\mathrm{tr}(\bm{\Omega}^{2}),\quad\mathcal{I}_{3}=\mathrm{tr}(\bm{S}^{3}),\quad\mathcal{I}_{4}=\mathrm{tr}(\bm{\Omega}^{2}\bm{S}),\quad\mathcal{I}_{5}=\mathrm{tr}(\bm{\Omega}^{2}\bm{S}^{2}),$ (16)
and $\bm{\mathcal{T}}^{(k)}$ are the symmetric tensor basis functions (the complete set is listed in Table 1). The coefficients $G^{(i)}$ are scalar, non-linear functions which depend on the five invariants $\mathcal{I}_{1}...\mathcal{I}_{5}$ and must be determined. If $G^{(1)}=-0.09,G^{(n)}=0$, the NLEVM degenerates to the classical $k-\epsilon$ model. When the NLEVM was proposed, it was impossible to find good approximations for these functions and as a result, it did not receive adequate attention. This hurdle however was overcome with the help of machine learning [15] where $G^{(i)}$ were learned from high-fidelity simulation data. Neural networks with parameters $\bm{\theta}$ were employed for the coefficients i.e. $G_{\bm{\theta}}^{(i)}$ and:
$\bm{b}_{\bm{\theta}}=\sum_{i=1}^{10}G_{\bm{\theta}}^{(i)}(\underbrace{\mathcal{I}_{1}...\mathcal{I}_{5}}_{\begin{subarray}{c}Scalar\\ Invariants\end{subarray}})\bm{\mathcal{T}}^{(i)};\quad\bm{\tau}_{\bm{\theta}}=2k\bm{b}_{\bm{\theta}}+\frac{2k}{3}\bm{I}.$ (17)
@@
As in [22], we employ the following prior for the NN parameters $\bm{\theta}$:
$p(\bm{\theta}\mid\nu)=\mathcal{N}(\bm{\theta}|0,\nu^{-1}\bm{I}_{d_{\theta}}),\quad p(\nu)=Gamma(\nu|a_{0},b_{0}),$ (18)
where $d_{\theta}=dim(\bm{\theta})$ and a Gamma hyperprior was used for the common precision hyperparameter $\nu$ with $(a_{0},b_{0})=(1.0,0.02)$. The resulting prior has the density of a Student’s $\mathcal{T}$-distribution centered at zero, which is obtained by analytically marginalizing over the hyperparameter $\nu$ [68].
Similarly however to the most widely used RANS closure models, such as the Launder-Sharma $k-\epsilon$ [56] or Wilcox’s $k-\omega$ [57], which are based on the Boussinesq turbulent-viscosity hypothesis, the present model of $\bm{\tau}_{\bm{\theta}}$ postulates that the RS tensor at each point in the problem domain depends on the flow features at the same point (locality assumption). This is a very strong assumption for flows that exhibit strong inhomogeneity [2].
$\bm{\mathcal{T}}^{(1)}$ = $\bm{S}$, $\bm{\mathcal{T}}^{(6)}$ = $\bm{\Omega}^{2}\bm{S}+\bm{S}\bm{\Omega}^{2}-\frac{2}{3}\mathrm{tr}(\bm{S}\bm{\Omega}^{2})\bm{I}$,
$\bm{\mathcal{T}}^{(2)}$ = $\bm{S}\bm{\Omega}-\bm{\Omega}\bm{S}$, $\bm{\mathcal{T}}^{(7)}$ = $\bm{\Omega}\bm{S}\bm{\Omega}^{2}+\bm{\Omega}^{2}\bm{S}\bm{\Omega}$,
$\bm{\mathcal{T}}^{(3)}$ = $\bm{S}^{2}-\frac{1}{3}\mathrm{tr}(\bm{S}^{2})\bm{I}$, $\bm{\mathcal{T}}^{(8)}$ = $\bm{S}\bm{\Omega}\bm{S}^{2}-\bm{S}^{2}\bm{\Omega}\bm{S}$,
$\bm{\mathcal{T}}^{(4)}$ = $\bm{\Omega}^{2}-\frac{1}{3}\mathrm{tr}(\bm{\Omega}^{2})\bm{I}$, $\bm{\mathcal{T}}^{(9)}$ = $\bm{\Omega}^{2}\bm{S}^{2}+\bm{S}^{2}\bm{\Omega}^{2}-\frac{2}{3}\mathrm{tr}(\bm{S}^{2}\bm{\Omega}^{2})\bm{I}$,
$\bm{\mathcal{T}}^{(5)}$ = $\bm{\Omega}\bm{S}^{2}-\bm{S}^{2}\bm{\Omega}$, $\bm{\mathcal{T}}^{(10)}$ = $\bm{\Omega}\bm{S}^{2}\bm{\Omega}^{2}-\bm{\Omega}^{2}\bm{S}^{2}\bm{\Omega}$,
@@ 2.2.2 Stochastic discrepancy term to RS
$\bm{\epsilon}_{\tau}=\bm{W}\bm{E_{\tau}}.$ (19)
It is based on considering $N_{d}$ subdomains of the problem domain and assuming that for all grid points in a certain subdomain, the corresponding RS discrepancy terms are identical. This can be expressed as in Equation (19) above where the entries of $\bm{W}$ are $1$ if a corresponding grid point (row of $\bm{W}$) belongs in a certain subdomain (column of $\bm{W}$) and 00 otherwise. A non-binary $\bm{W}$ would also be possible, although in this case its physical interpretation in terms of subdomains would be occluded. The vector $\bm{E}_{\tau}=\{\bm{E}_{\tau,J}\}_{J=1}^{N_{d}}$ contains therefore the RS discrepancy terms for each subdomain, e.g. $dim(\bm{E}_{\tau,J})=3$ for a two-dimensional flow. The dimensionality reduction scheme resembles Principal Component Analysis (PCA) [69]. In contrast to the latter however, we never have data/observations of the vectors to be reduced, i.e. $\bm{\epsilon_{\tau}}$ in our case. These are inferred implicitly from the LES data and simultaneously with $\bm{W}$. In the ensuing numerical illustrations, the division into subdomains is done in a regular manner and $\bm{W}$ is prescribed a priori. One could nevertheless readily envision a learnable $\bm{W}$ or even an adaptive refinement into subdomains.
The incorporation of the model discrepancy variables $\bm{\epsilon}_{\tau}$ or $\bm{E}_{\tau}$ at each grid-point or subdomain introduces redundancies i.e. there would be an infinity of combinations of $\bm{\theta}$ and $\bm{\epsilon}_{\tau}$/$\bm{E}_{\tau}$ that could fit the data equally well. In order to address this issue, we invoke the concept of sparsity which has been employed in similar situations in the context of physical modeling [70, 71]. To this end, we make use of a sparsity-enforcing Bayesian prior based on the Automatic Relevance Determination (ARD) [72, 73]. In particular:
$p(\bm{E}_{\tau}|\bm{\Lambda})=\prod_{J=1}^{N_{d}}p(\bm{E}_{\tau,J}|\bm{\Lambda}^{(J)})=\prod_{J=1}^{N_{d}}\mathcal{N}\left(\bm{E}_{\tau,J}~{}|\bm{0},~{}diag(\bm{\Lambda}_{J})^{-1})\right),$ (20)
@@
$p(\bm{\Lambda})=\prod_{J=1}^{N_{d}}\prod_{\ell=1}^{L}Gamma(\Lambda_{J,\ell}~{}|\alpha_{0},\beta_{0}),$ (21)
where $\Lambda_{J,\ell}$ denotes the $\ell^{th}$ entry (e.g. $L=3$ for two-dimensional flows) of the vector of precision hyperparameters in subdomain $J$. We note that when $\Lambda_{J,\ell}\to\infty$, then the corresponding model discrepancy term $E_{\tau,J,\ell}\to 0$. The resulting prior for $\bm{E}_{\tau}$ arising by marginalizing the hyperparameters $\bm{\Lambda}$ is a light-tailed, Student’s t-distribution [74] that promotes solutions in the vicinity of 00 unless strong evidence in the data suggests otherwise. The hyperparameters $\alpha_{0},\beta_{0}$ are effectively the only ones that need to be provided by the analyst. We advocate very small values ($\alpha_{0}=\beta_{0}=10^{-3}$ was used in the ensuing numerical illustrations) which correspond to an uninformative prior [75]. This is consistent with the values adopted in the original paper where the ARD prior was employed in conjunction with variational inference i.e. in [75].
The dimension of the $\bm{\epsilon}_{\tau}$ is high i.e. $3N$ in 2D (and $6N$ in 3D) where $N$ is the total number of grid points. In contrast, the dimension of $\bm{\theta}$ is independent of the grid as the parametric part of the closure model is the same at all grid points. Depending on the amount of data, one can envision cases where multiple combinations of $\bm{\epsilon}_{\tau}$ and $\bm{\theta}$ could fit the data equally well. Non-zero values of $\bm{\epsilon}_{\tau}$ would imply model errors at the corresponding grid points. Ceteris paribus, we would prefer the combination with the least number of model errors or equivalently the one for which our parametric model can capture for most of the domain the closure accurately. It is indeed these types of solutions/combinations that the ARD prior promotes.